IF YOU HAVE TWO TRIANGLES AND THEY ARE SIMILAR TO EAC OTHER, ONE NAMED PQR AND THE OTHER NAMED XYZ, AND PQR SIDES ARE LABELED SIDE QP=9 AND SIDE PR=27. AND TRIANGLE XYZ HAS YX=5. AND QP=YX AND PR=ZX. AND I HAD TO FIND THE MISSING SIDE WHICH IS ZX, WHAT WILL BE THE ANSWER.

THANKS,
MIKAYLA
P.S. IF YOU DON'T KNOW THE ANSWER, THANK YOU ANYWAY

Answer by RAY100(1637)

a rough sketch is important, size is not
.
Label carefully, p matches x, q matches y, and r matches z
.
pq and xy are proportional as triangles are similar,,,SF = 5/9,,,,,,, (scale factor)
.
applying scale factor to pr =27,,,,27 *5/9 = 15 for XZ
.
applying scale factor to zy =5,,,,, 5*9/5 = 9 for RQ
.
Thank You and have a good day